OpenStudy (anonymous):

Right, here we go: z = -3+4i, is one root of the polynomial... \[z^4 + 2z^3 + 9z^2 -52z +200 =0\] Find the three remaining roots.

OpenStudy (anonymous):

z1=-3+4i, z2=-3-4i, and then divide the equation by (z-z1)(z-z2) you will get z3=2-2i,z4=2+2i

OpenStudy (anonymous):

actually it is (z^2+6z+25)(z^2-4z+8)=0